OpenStudy (anonymous):

Lex f(x) =5x ^{6}sqrt{x}+6/(x ^{3}sqrt{x}). Find f'(x)

OpenStudy (anonymous):

is this what you mean \[5\,{x}^{13/2}+6\,{x}^{-7/2}\]

OpenStudy (anonymous):

\[5x ^{6}\sqrt{x}+6/(x ^{3}\sqrt{x})\]

OpenStudy (anonymous):

lol same thing :)

OpenStudy (anonymous):

whoops, then yes lol

OpenStudy (anonymous):

well use the same rule i showed you earlier plus this rule f'(x)=a'(x)+b'(x) something like that meaning you can take the derivitive of each term individually.

OpenStudy (anonymous):

So then is it \[65/2x ^{11/2}-21x ^{-9/2}\]

OpenStudy (anonymous):

x^(13/2)+6/x^(7/2) 13/2 x^2(13/2-1) + -7/2*6^(-7/2-1)

OpenStudy (anonymous):

after that just multiply out whatever needs to be multiplied the (-7/2) * 6 subtract the ones inside the exponents to make the fractions simplified then result is (65/2)*x^(11/2)-21/x^(9/2) \[{\frac {65}{2}}\,{x}^{11/2}-21\,{x}^{-9/2}\]

OpenStudy (anonymous):

You are so right :D wow you dont need me at all lol.

OpenStudy (anonymous):

Haha thanks! so what is f'(6)?

OpenStudy (anonymous):

That answer above was wrong

OpenStudy (anonymous):

plug in 6 at all the xs (655050233/2592)*sqrt(6)

OpenStudy (anonymous):

which is 619035.041181

OpenStudy (anonymous):

they're both incorrect

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I'm not sure why.

OpenStudy (anonymous):

i see i think it is because of the 21

OpenStudy (anonymous):

how odd i redid all and still get same :(

OpenStudy (anonymous):

and i did it with my calculator and it got me same result manually

OpenStudy (anonymous):

hmmm....so is the answer still the same?

OpenStudy (anonymous):

you get 619035.041181 from the 1st term. Did you subtract the 2nd?

OpenStudy (anonymous):

I typed that in as well and those aren't correct.

OpenStudy (anonymous):

Mb limit for decimals or asked for f(6), not f'(6)?

OpenStudy (anonymous):

f'(x)=2/3[x^(15/2)]-12/5[x^(-5/2)] + c......(X) (c is a constant) put x=6 and get c the resubstitute in eq. X and get f'(x)......